OpenStudy (anonymous):

help please ( 3.26x 106-3 mg ) - ( 7.88x 10^-5 µg )

OpenStudy (anonymous):

1 miligram is the same as 10^-3 grams and a micro gram is the same as 10^-6

OpenStudy (anonymous):

yes , i know , but the last answer is in gram , miligram or microgram ?

OpenStudy (anonymous):

At the end you will get an answer in grams

OpenStudy (anonymous):

okay . can you show me your calculations so that I can compare it with mine

OpenStudy (anonymous):

first what do you mean by "106-3 mg"

OpenStudy (anonymous):

nvm 10^-6?

OpenStudy (anonymous):

act it is 3.26 x 10 ^ -3 . typing error

OpenStudy (anonymous):

3.26*10^-3*10^-3=3.26*10^-6 7.88*10^-5*10^-6=7.88*10^-11 put that into a calculator:you get 0.00000325992 grams

OpenStudy (anonymous):

convert to 3 sig-figs: 3.25*10^6grams

OpenStudy (anonymous):

you can change that to micro grams if you want making in 3.25 ugrams

OpenStudy (anonymous):

okay . thank you so much

OpenStudy (anonymous):

np